# Inverse source problem with a posteriori interior measurements for space-time fractional diffusion equations

### v1

@@ Abstract
This paper investigates an inverse source problem for space-time fractional diffusion equations from a posteriori interior measurements. The uniqueness result is established by the memory effect of fractional derivatives and the unique continuation property. For the numerical reconstruction, the inverse problem is reformulated as an optimization problem with the Tikhonov regularization. We use the Levenberg-Marquardt method to identity the unknown source from noisy measurements. Finally, we give some numerical examples to illustrate the efficiency and accuracy of the proposed algorithm.
@@ 1 Introduction and the main result
Along with the intensive attention paid to forward problems for space-time fractional diffusion equations, there has been a rapid growth in research on the corresponding inverse problems. For example, Zhang et al. 2018Bayesian employed a Bayesian approach to address the inverse problem for a time-space fractional diffusion equation, but they did not give a uniqueness result. Li and Wei 2018An investigated the identification of a time-dependent source component using initial and boundary data as well as extra measurement data at an inner location in a space-time fractional diffusion problem. Janno 2020Determination investigated an inverse problem for identifying time-dependent sources and parameters in nonlocal diffusion and wave equations from final data. Li et al. 2020Tikhonov employed a Tikhonov regularization approach to solve the inverse problem of identifying a space-dependent source term in a time-space fractional diffusion equation. Tatar et al. in 2015Determination ; 2016Simultaneous ; 2015An solved an inverse space-dependent source problem and identified the orders of time and space fractional derivatives for a time-space fractional diffusion equation. Jia et al. in 2017Harnack's investigated Harnack’s inequality for a space-time fractional diffusion equation and applications to an inverse source problem. In 2021A fractional , Djennadi et al. studied inverse backward and source problems in the time-space-fractional diffusion equations by fractional Tikhonov regularization method. Nie et al. in 2023An inverse investigated an inverse random source problem for the time-space fractional diffusion equation driven by fractional Brownian motion. Van et al. used the quasi-reversibility method to study an inverse source problem for time-space fractional parabolic equations in 2023The .
@@ 2 Preliminaries
[…]
@@
$\mathcal{L}[E_{\alpha,1}(-\lambda t^{\alpha})](z)=\frac{z^{\alpha-1}}{\lambda+z^{\alpha}},~{}~{}z\in\mathbb{C},~{}{\alpha>0}.$
@@
For the definition of the fractional elliptic operator, we first introduce the eigensystem $\{(\lambda_{n},\varphi_{n})\}_{n=1}^{\infty}$ of $\mathcal{A}$, namely $\mathcal{A}\varphi_{n}=\lambda_{n}\varphi_{n}$ with $\varphi_{n}\in H^{2}(\Omega)\cap H_{0}^{1}(\Omega)$. Then it is known that $0<\lambda_{1}\leq\lambda_{2}\leq\cdots$, $\lambda_{n}\rightarrow\infty$ as $n\to\infty$, and $\{\varphi_{n}\}$ forms a complete orthonormal basis of $L^{2}(\Omega)$. Now we can define the fractional Sobolev space $D(\mathcal{A}^{\beta})$ and the corresponding fractional elliptic operator $\mathcal{A}^{\beta}$ as
$D(\mathcal{A}^{\beta})$ :=$:=\left\{u\in L^{2}(\Omega)\mid\|u\|_{D(\mathcal{A}^{\beta})}^{2}:=\sum_{n=1}^{\infty}\left|\lambda_{n}^{\beta}(u,\varphi_{n})\right|^{2}<\infty\right\},$
@@
[…]
$=C\|\mu\|^{2}_{L^{2}(0,T)}\sum_{n=1}^{\infty}\left(1-E_{\alpha,1}(-\lambda_{n}^{\beta/2}T^{\alpha})\right)^{2}f_{n}^{2}$
@@
Moreover, we can analytically continue both sides of (3.1), so that it holds for $\eta\in\mathbb{C}\setminus\{-\lambda^{\beta}_{n}\}_{n=1}^{\infty}$. Then for any $n=1,2,\dots$, we can take a sufficiently small circle centered at $-\lambda^{\beta}_{n}$ excluding distinct eigenvalues and integrate (3.1) on this circle to obtain
@@
Simultaneously, it is readily seen that $u_{n}$ satisfies the elliptic equation $(\mathcal{A}-\lambda_{n})u_{n}=0$ in $\Omega$. Then it follows from the unique continuation for elliptic equations that $u_{n}\equiv 0$ in $\Omega$ for all $n=1,2,\dots$. Together with the complete orthogonality of $\{\varphi_{n}\}$ in $L^{2}(\Omega)$, this implies $\phi=0$ in $\Omega$ as the initial value of (1.1). Consequently, the uniqueness for the forward problem guarantees $u\equiv 0$ in $\Omega\times(0,T)$, which completes the proof. ∎
@@ 4 Levenberg-Marquardt method
and two vectors $\bm{a}^{k+1}=(a_{1}^{k+1},\dots,a_{N}^{k+1}),\bm{a}^{k}=(a_{1}^{k},\dots,a_{N}^{k})\in\mathbb{R}^{N}$. We identify approximations $f^{k+1},f^{k}\in\Phi^{N}$ with vectors $a^{k+1},a^{k}\in\mathbb{R}^{N}$, respectively.
@@
$W^{k}=\left(\left(u^{\delta}-u(\,\cdot\,,\,\cdot\,;\bm{a}^{k}),\frac{\partial}{\partial a_{i}^{k}}u(\,\cdot\,,\,\cdot\,;\bm{a}^{k})\right)_{L^{2}(\omega\times(T_{1},T))}\right)_{N\times 1}.$
@@ 5 Numerical experiments
$u^{\delta}(x,t)=u(x,t;f_{\mathrm{true}})+\epsilon\,u(x,t;f_{\mathrm{true}})(2\,\mathrm{rand}(-1,1)-1),\quad(x,t)\in\omega\times(T_{1},T),$
where $\mathrm{rand}(-1,1)$ denotes a random number distributed uniformly in $(-1,1)$. The corresponding noise level is then calculated by $\delta=\|u^{\delta}-u(\,\cdot\,,\,\cdot\,;f_{\mathrm{true}})\|_{L^{2}(\omega\times(T_{1},T))}$, where we recall that $\omega$ is a nonempty subdomain of $\Omega$. To evaluate the accuracy of the numerical solution, we compute the relative error defined by
@@
where $\eta>1$ is a constant usually taken heuristically to be 1.01. If the noise level is 00, then we take $K=40$ for the following examples. The regularization parameter $\rho_{k}$ at the $k$th step is selected as the sigmoid-type function based on its properties given by
@@
Reconstruction results of the source term in $\Omega\setminus\omega$ with various noise levels are shown in Figures 12. We can observe that our algorithm perform worse as $T_{1}$ increases. This result is definitely consistent with our expectation because the reduction in available information makes the inversion of the algorithm more difficult. We can see from Figures 12 and Tables 12 that the numerical results match the true solution rather satisfactory up to $1\%$ noise added in the noiseless data for different choices of $\alpha$ and $\beta$. Finally, based on the analysis of numerical results, we verify the validity of the theory and the stability of the algorithm.
@@ 6 Conclusion
We recover the space-dependent source term from a posteriori interior measurements for the space-time fractional diffusion equations in this paper. The uniqueness of the corresponding inverse problem was established by employing memory effect of fractional derivative, unique continuation property and Duhamel’s principle for space-time fractional diffusion equations. We use the Levenberg-Marquardt method to solve the inverse problem. Finally, numerical results verify the correctness of the proposed theory.
@@ Acknowledgments
The second author is supported by National Natural Science Foundation of China (no. 12271277) and Ningbo Youth Leading Talent Project (no. 2024QL045). The third author is supported by JSPS KAKENHI Grant Numbers JP22K13954, JP23KK0049 and Guangdong Basic and Applied Basic Research Foundation (No. 2025A1515012248). This work is partly supported by the Open Research Fund of Key Laboratory of Nonlinear Analysis & Applications (Central China Normal University), Ministry of Education, China.

### v2

@@ Abstract
This paper investigates an inverse source problem for space-time fractional diffusion equations from a posteriori interior measurements. The uniqueness result is established by the memory effect of fractional derivatives and the unique continuation property. For the numerical reconstruction, the inverse problem is reformulated as an optimization problem with the Tikhonov regularization. We use the Levenberg-Marquardt method to identify the unknown source from noisy measurements. Finally, we give some numerical examples to illustrate the efficiency and accuracy of the proposed algorithm.
@@ 1 Introduction and the main result
Along with the intensive attention paid to forward problems for space-time fractional diffusion equations, there has been a rapid growth in research on the corresponding inverse problems. For example, Zhang et al. 2018Bayesian employed a Bayesian approach to address the inverse problem for a time-space fractional diffusion equation, but they did not give a uniqueness result. Li and Wei 2018An investigated the identification of a time-dependent source component using initial and boundary data as well as extra measurement data at an inner location in a space-time fractional diffusion problem. Janno 2020Determination investigated an inverse problem for identifying time-dependent sources and parameters in nonlocal diffusion and wave equations from final data. Li et al. 2020Tikhonov employed a Tikhonov regularization approach to solve the inverse problem of identifying a space-dependent source term in a time-space fractional diffusion equation. Tatar et al. in 2015Determination ; 2016Simultaneous ; 2015An solved an inverse space-dependent source problem and identified the orders of time and space fractional derivatives for a time-space fractional diffusion equation. Jia et al. in 2017Harnack's investigated Harnack’s inequality for a space-time fractional diffusion equation and applications to an inverse source problem. In 2021A fractional , Djennadi et al. studied backward and inverse source problems in the time-space-fractional diffusion equations by fractional Tikhonov regularization method. Nie et al. in 2023An inverse investigated an inverse random source problem for the time-space fractional diffusion equation driven by fractional Brownian motion. Van et al. used the quasi-reversibility method to study an inverse source problem for time-space fractional parabolic equations in 2023The .
@@ 2 Preliminaries
In this section, we first fix notations and introduce tools for later use, and then mention the well-posedness of the forward problem. Let $L^{2}(\Omega)$ be a usual $L^{2}$-space with the inner product (,)($\cdot\,$$\cdot\,$), and $H_{0}^{1}(\Omega),H^{2}(\Omega)$ etc. denote the usual Sobolev spaces. By $H^{\alpha}(0,T)$, we denote the Sobolev-Slobodeckij space (e.g., Adams AF03 ).
@@
$\mathcal{L}\left\{t^{\alpha-1}E_{\alpha,\beta}\left(\pm\lambda t^{\alpha}\right)\right\}(z)=\frac{z^{\alpha-\beta}}{z^{\alpha}\mp\lambda},~~\Re(z)>|\lambda|^{\frac{1}{\alpha}},~~\alpha,\beta>0.$
@@
For the definition of the fractional elliptic operator, we first introduce the eigensystem $\{(\lambda_{n},\varphi_{n})\}_{n=1}^{\infty}$ of $\mathcal{A}$, namely $\mathcal{A}\varphi_{n}=\lambda_{n}\varphi_{n}$ with $\varphi_{n}\in H^{2}(\Omega)\cap H_{0}^{1}(\Omega)$. Then it is known that $0<\lambda_{1}\leq\lambda_{2}\leq\cdots$, $\lambda_{n}\rightarrow\infty$ as $n\to\infty$, and $\{\varphi_{n}\}_{n=1}^{\infty}$ form a complete orthonormal basis of $L^{2}(\Omega)$. Now we can define the fractional Sobolev space $D(\mathcal{A}^{\beta})$ and the corresponding fractional elliptic operator $\mathcal{A}^{\beta}$ as
$D(\mathcal{A}^{\beta})$ $:=\left\{u\in L^{2}(\Omega)\mid\|u\|_{D(\mathcal{A}^{\beta})}^{2}:=\sum_{n=1}^{\infty}\left|\lambda_{n}^{\beta}(u,\varphi_{n})\right|^{2}<\infty\right\},$
@@
$\leq\|\mu\|^{2}_{L^{2}(0,T)}\sum_{n=1}^{\infty}\left(\lambda_{n}^{\beta/2}\int_{0}^{T}t^{\alpha-1}E_{\alpha,\alpha}(-\lambda_{n}^{\beta}t^{\alpha})dt\right)^{2}f_{n}^{2}$
$\leq C\|\mu\|^{2}_{L^{2}(0,T)}\sum_{n=1}^{\infty}\left(1-E_{\alpha,1}(-\lambda_{n}^{\beta/2}T^{\alpha})\right)^{2}f_{n}^{2}$
@@
Moreover, we can analytically continue both sides of (3.1), so that it holds for $\eta\in\mathbb{C}\setminus\{-\lambda^{\beta}_{n}\}_{n=1}^{\infty}$. Then for any $n=1,2,\dots$, we can take a sufficiently small circle centered at $\lambda^{\beta}_{n}$ excluding distinct eigenvalues and integrate (3.1) on this circle to obtain
@@
Simultaneously, it is readily seen that $u_{n}$ satisfies the elliptic equation $(\mathcal{A}-\lambda_{n})u_{n}=0$ in $\Omega$. Then it follows from the unique continuation for elliptic equations that $u_{n}\equiv 0$ in $\Omega$ for all $n=1,2,\dots$. Together with the complete orthogonality of $\{\varphi_{n}\}_{n=1}^{\infty}$ in $L^{2}(\Omega)$, this implies $\phi=0$ in $\Omega$ as the initial value of (1.1). Consequently, the uniqueness for the forward problem guarantees $u\equiv 0$ in $\Omega\times(0,T)$, which completes the proof. ∎
@@ 4 Levenberg-Marquardt method
and two vectors $\bm{a}^{k+1}=(a_{1}^{k+1},\dots,a_{N}^{k+1}),\bm{a}^{k}=(a_{1}^{k},\dots,a_{N}^{k})\in\mathbb{R}^{N}$. We identify approximations $f^{k+1},f^{k}\in\Phi^{N}$ with vectors $\bm{a}^{k+1},\bm{a}^{k}\in\mathbb{R}^{N}$, respectively.
@@
$\bm{W}^{k}=\left(\left(u^{\delta}-u(\,\cdot\,,\,\cdot\,;\bm{a}^{k}),\frac{\partial}{\partial a_{i}^{k}}u(\,\cdot\,,\,\cdot\,;\bm{a}^{k})\right)_{L^{2}(\omega\times(T_{1},T))}\right)_{N\times 1}.$
@@ 5 Numerical experiments
$u^{\delta}(x,t)=(1+\epsilon\,\mathrm{rand}(-1,1))u(x,t;f_{\mathrm{true}}),\quad(x,t)\in\omega\times(T_{1},T),$
where $\mathrm{rand}(-1,1)$ denotes a random number distributed uniformly in (1,1)(-1,1). The corresponding noise level is then calculated by $\delta=\|u^{\delta}-u(\,\cdot\,,\,\cdot\,;f_{\mathrm{true}})\|_{L^{2}(\omega\times(T_{1},T))}$, where we recall that $\omega$ is a nonempty subdomain of $\Omega$. To evaluate the accuracy of the numerical solution, we compute the relative error defined by
@@
where $\eta>1$ is a constant usually taken heuristically to be 1.01. If the noise level is 0, then we take $K=40$ for the following examples. The regularization parameter $\rho_{k}$ at the $k$th step is selected as the sigmoid-type function based on its properties given by
@@
Reconstruction results of the source term in $\Omega\setminus\omega$ with various noise levels are shown in Figures 12. We can observe that our algorithm performs worse as $T_{1}$ increases. This result is definitely consistent with our expectation because the reduction in available information makes the inversion of the algorithm more difficult. We can see from Figures 12 and Tables 12 that the numerical results match the true solution rather satisfactory up to $1\%$ noise added in the noiseless data for different choices of $\alpha$ and $\beta$. Finally, based on the analysis of numerical results, we verify the validity of the theory and the stability of the algorithm.
@@ 6 Conclusion
We recover the space-dependent source term from a posteriori interior measurements for the space-time fractional diffusion equations in this paper. The uniqueness of the corresponding inverse problem is established by employing memory effect of fractional derivative, unique continuation property and Duhamel’s principle for space-time fractional diffusion equations. We use the Levenberg-Marquardt method to solve the inverse problem. Finally, numerical results verify the correctness of the proposed theory.
@@ Acknowledgments
The authors thank the anonymous referees for valuable comments. The second author is supported by National Natural Science Foundation of China (no. 12271277) and Ningbo Youth Leading Talent Project (no. 2024QL045). The third author is supported by JSPS KAKENHI Grant Numbers JP22K13954, JP23KK0049 and Guangdong Basic and Applied Basic Research Foundation (No. 2025A1515012248). This work is partly supported by the Open Research Fund of Key Laboratory of Nonlinear Analysis & Applications (Central China Normal University), Ministry of Education, China.